OpenStudy (anonymous):

The equation of an ellipse is shown. (x+2)^2/100 + (y+7)^2/36 = 1 What are the coordinates of the foci?

OpenStudy (anonymous):

(-10, -7) and (6, -7) (-6, 7) and (10, 7) (-2, -15) and (-2, 1) (2, -1) and (2, 15)

OpenStudy (anonymous):

I think it is (-6, 7) and (10, 7)

OpenStudy (jdoe0001):

hmm how did you get (-6, 7) and (10, 7) ?

OpenStudy (mathmate):

The translation will be more obvious if you write it as: \[\frac{(x-(-2))^2}{10^2}+\frac{(y-(-7))^2}{6^2}\]

OpenStudy (mathmate):

In general, the centre of an ellipse is translated from the origin (0,0) to (h,k) using: \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}\] where a, b are the horizontal and vertical axes. and the foci are (h+c, k), (h-c,k) when a>b, (h,k-c), (h,k+c) when a<b. also, 2c is the focal distance, where c^2=|a^2-b^2|